OpenStudy (anonymous):

find the coordinare of the midpoint of segment LS given L(-8,8) and S(-4,4)

OpenStudy (anonymous):

MP Formula : ( ((x1 + x2)/2), ((y1 + y2)/2) ) -8 + -4 = -12 , -12/2 = -6, 8+4 = 12, 12/2 = 6 so (-6,6)

OpenStudy (mertsj):